OpenStudy (anonymous):

Solve the absolute value equation |2n - 4| = 8. One solution is n = 6, what is the other solution?

OpenStudy (anonymous):

|2n - 4| = 8 2n - 4 = 8 or 2n - 4 = -8 try solving them

hero (hero):

|2n - 4| = 8 2n - 4 = ±8

OpenStudy (anonymous):

thank u